OpenStudy (anonymous):

A shirt cost $9.88 What is the price of the item after a 10% increase with sales tax? (Sales Tax: 8.25%)

OpenStudy (anonymous):

So, lets try it this way. Tax is added to the final price. So lets first deal with the 10% Increase. So first, know that 0.1 = 10% 1.) Then: $9.88 * (0.1) = .988 which is rounded to $0.99 So then you add the shirt price plus the increase: 9.88 + .99 But before we add it, we then do the tax. The tax is 8.25% = 0.825 2) So repeat the process of part 1: 9.88 * 0.0825 = about .82 3) add the parts now: $9.88 + (Increase) $0.99 + (tax) $0.82 = $11.69 Hope this helps

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Welcome :)